# Key Issues Review: Useful autonomous quantum machines

### v2

@@ Abstract
Controlled quantum machines have matured significantly. A natural next step is to increasingly grant them autonomy, freeing them from time-dependent external control. For example, autonomy could pare down the classical control wires that heat and decohere quantum circuits; and an autonomous quantum refrigerator recently reset superconducting qubits to near their ground states, as is necessary before a computation. Which fundamental conditions are necessary for realizing useful autonomous quantum machines? Inspired by recent quantum thermodynamics and chemistry, we posit conditions analogous to DiVincenzo’s criteria for quantum computing. Furthermore, we illustrate the criteria with multiple autonomous quantum machines (refrigerators, circuits, clocks, etc.) and multiple candidate platforms (neutral atoms, molecules, superconducting qubits, etc.). Our criteria are intended to foment and guide the development of useful autonomous quantum machines.
@@
    Autonomous quantum machines of a third type could detect, transduce, and transmit energy. Nature has produced machines of this type. Photoisomers are molecules operative across nature and technologies [98, 99, 100, 101]. Examples include retinal, found in the protein rhodopsin in our eyes [102]. A photoisomer has one shape in thermal equilibrium at biological temperatures. Absorbing a photon—say, from the sun—enables the molecule to switch configurations. The switch can galvanize chemical reactions leading to the experience of sight. Photoisomers exhibit quantum phenomena including coherence relative to the energy eigenbasis, as well as nonadiabatic evolution. Beyond photoisomers, chemistry features other autonomous quantum machines: Photosynthetic complexes transform light into chemical energy [103], enzymes interconvert molecules with help from tunneling [104], etc.
@@ I Definitions
Before presenting our criteria, we specify meanings for machine, autonomous, and quantum. Alternative definitions may exist, as the terms are broad. Still, we believe our definitions to be reasonable. Specifying them will clarify our criteria:
@@
    Autonomous: A machine is autonomous if its total microscopic Hamiltonian is not changed by any agent using the machine. An autonomous machine’s user performs no thermodynamic work on the machine during the machine’s operation. Work tends to be defined through changes in the system-of-interest Hamiltonian, in quantum thermodynamics [80].555 For example, an adenosine-triphosphate (ATP) molecule provides chemical work via hydrolyzation: one of its phosphates splits off from the rest of the molecule, which becomes adenosine diphosphate (ADP). The cleaving of the bond releases energy and changes the molecule’s Hamiltonian.
@@
    Relatedly, this Perspective does not posit criteria under which autonomous quantum machines generally outperform all classical counterparts. Rather, the Perspective provides criteria under which autonomous quantum machines are useful. We believe this goal to be worthwhile, ambitious, and reasonable for the near future. Farther in the future, another Perspective may posit criteria under which autonomous quantum machines generally outperform all classical counterparts.
@@ A Access to useful energy
Other energy sources provide work indirectly. They are nonequilibrium systems that contain free energy (or a generalization thereof), which a machine can harvest to produce work. Often, such systems have gradients of temperature, chemical potential, or other generalized thermodynamic forces. The commonest example consists of two heat baths. One, at a temperature $T_{\mathcal{C}}$, is out of equilibrium with a heat bath at a temperature $T_{\mathcal{H}}>T_{\mathcal{C}}$. Heat flows between the baths. This heat is not work, so the temperature gradient does not provide work directly. However, a machine can siphon off a fraction of the current and transform it partially (in accordance with Carnot’s theorem) into work. Such machines are called absorption machines [74]. Superconducting qubits have been fueled by heat baths of two types: resonators [122] and thermal fields propagating through microwave waveguides [75]. Despite providing energy, baths can threaten the purity of a machine’s state. Yet, as discussed under criterion F, non-Markovian baths, which retain memories, can revive lost purity.
@@ B Processing unit or target
A machine’s purpose can be to operate on a target. For instance, a refrigerator’s target is the system that undergoes cooling. An engine’s target may be a battery that stores the energy extracted by the engine. A drone’s target is the package to be delivered.
@@ C Interactions among the machine’s components
Figure 2: Interactions amongst components: The quantum absorption refrigerator and clock evolve under similar Hamiltonians. However, the machines’ qudits have different energy gaps: the refrigerator’s $\Delta_{\mathcal{H}}<\Delta_{\mathcal{C}}$, whereas the clock’s $\Delta_{\mathcal{H}}>\Delta_{\mathcal{C}}$. Also, the number of the target’s energy gaps can differ between machines. The gaps, with the baths’ temperatures ($T_{\mathcal{H}}$ and $T_{\mathcal{C}}$), determine the directions in which energy flows.
@@
Effective three-body interactions are necessary for autonomous cooling [25]. They can be effected perturbatively with simultaneous two-body interactions [32, 34, 46, 75]. The exchange (1) occurs only if $\Delta_{\mathcal{H}}+\Delta_{\mathcal{T}}=\Delta_{\mathcal{C}}$—under a condition called strict, or microscopic, energy conservation [131, 132]. The thermal qubits’ gaps and temperatures bias the interaction (1) to the right. Ending in its ground state, $\mathcal{T}$ has undergone cooling by the refrigerator.
@@
Identity operators $\mathbbm{1}$ are tensored on wherever necessary for each operator to act on the full Hilbert space.
@@ D Timekeeping mechanism
Timekeeping devices include the switches introduced in the previous subsection; clocks, which tick regularly; and timers, which announce when programmable amounts of time have passed. Below, we review a simple autonomous quantum clock. Then, we discuss timekeeping mechanisms not formed from physical devices. Finally, we delineate three purposes of autonomous quantum machines’ timekeeping mechanisms.
@@ E Structural integrity
Yet a machine need not consist of physically separated components. A photoisomer functions as an autonomous quantum energy transmitter, as explained in the introduction. The molecule contains an autonomous quantum clock, according to [134]. Similarly, photosynthetic complexes [103] and enzymes [104] are autonomous machines modeled usefully with quantum theory. A molecule can therefore form an autonomous quantum machine. So can an atom, arguably.777 The atom interacts with heat baths whose frequencies are filtered. One might count the filters as parts of the machine. For example, three atomic energy levels can form an autonomous engine or refrigerator [2, 3, 25]. Not only natural particles, but also artificial devices can form self-contained autonomous quantum machines: Superconducting qubits cannot separate if printed on the same chip. Neither can the dopants that host qubits on solid-state surfaces; and neither can the gate electrodes that define quantum dots, being grown or printed on a semiconductor chip.
@@ G Output worth the input
Each figure of merit above measures an output’s quality. The efficiency and COP compare outputs with their input costs. Comprehensive figures of merit capture such comparisons [177]. Inputs can include energy, time, and control. We illustrate energy with a single-atom engine and a quantum refrigerator. Then, we distinguish two input costs: the fabrication of an autonomous quantum machine and the preparation of a fabricated machine for one trial. If the preparation is simple enough, and the machine executes enough trials, the total output can outweigh a large fabrication cost.
@@ I Mobility (optional)
    The signal must, given condition (1), have a sufficiently high probability of affecting $\mathcal{B}$. The excitation’s momentum should direct the signal toward $\mathcal{B}$, and no intervening medium should swallow the signal. Once the excitation arrives, $\mathcal{B}$ should have a high probability of absorbing it. Fermi’s golden rule mediates this probability, governing the rate $\Gamma_{\text{i}\mapsto\text{f}}$ at which $\mathcal{B}$ jumps from a state $\lvert\rm i\rangle$ to any energy-$E_{\rm f}$ state $\lvert\rm f\rangle$: $\Gamma_{\text{i}\mapsto\text{f}}=\frac{2\pi}{\hbar}\,|\langle\text{f}\rvert H^{\prime}\lvert\text{i}\rangle|^{2}\,\mu(E_{\rm f})$. $H^{\prime}$ denotes the excitation-induced perturbation to $\mathcal{B}$’s Hamiltonian. $\mu(E)$ denotes the density of $\mathcal{B}$’s energy-$E$ states. If $\mathcal{B}$ absorbs the signal through a photodetector, the external efficiency quantifies how effectively $\mathcal{B}$ satisfies requirement (2). One can achieve a high efficiency upon matching $\mathcal{A}$’s and $\mathcal{B}$’s impedances. Impedance matching has enabled nearly perfect photon absorption by a cavity-QED detector [184, 185]. The scheme leveraged the detector’s $\Lambda$-type energy-level structure, albeit nonautonomously. Apart from impedance matching, realistic interconnects entail other engineering challenges, such as losses in the interconnects.
@@
In another example, a qubit and a bosonic mode undergo a dispersive interaction. Such interactions are common in cavity QED, the qubit manifesting as a natural or artificial atom. The dispersive interaction is effective (approximate), and we omit the derivation [187]. During it, one attributes to the qubit an effective gap $\Delta$; to the mode an effective frequency $\omega$, a creation operator $a^{\dagger}$, and an annihilation operator $a$; and to the coupling a strength $\chi$. We set $\hbar=1$. The dispersive Hamiltonian is
@@
If the qubit is excited (in $\lvert 1\rangle$), it effectively adds $\chi$ to the mode’s frequency. Analogously, if the mode is occupied, it effectively adds $\chi$ to the qubit’s gap. Each subsystem therefore affects the other’s Hamiltonian.
@@ III Outlook
    Usable energy—especially energy extracted from heat baths—trades off with purity, as discussed below criterion F. Four strategies suggest themselves: (i) Take advantage of the “sufficient” in the “sufficient purity” criterion: Identify when purity is necessary, and waste no effort on maintaining unnecessary purity. (ii) Situate autonomous quantum machine in environments that stabilize desirable quantum states [67], using the toolkit of engineered dissipation [189]. The engineered may suggest that maintaining the environment requires work from an agent. This impression is misleading, however. For example, “the Zeno effect [ …] is largely passive” and can raise a system’s probability of remaining in its ground state [189]. Engineered dissipation generalizes non-Markovian environments, discussed in App. A. (iii) Resolving challenge 1 can resolve this tradeoff challenge.
@@ Acknowledgements.
N.Y.H. thanks Kenneth Brown for input about castle-in-the-air autonomous quantum machines, David Limmer for input about chemical machines, Noah Lupu-Gladstein for input about photodetectors, John Preskill for input about superconducting-qubit chips, and Nathan Schine for input about quantum-computing platforms. N.Y.H. and J.A.M.G. acknowledge support from the National Science Foundation (QLCI grant OMA-2120757), the John Templeton Foundation (award no. 62422), and NIST grant 70NANB21H055_0. P.E. and M.H. acknowledge support from the European Research Council (Consolidator grant “Co-coquest” 101043705), the European flagship on quantum technologies (“ASPECTS” consortium 101080167), and FQXi (FQXi-IAF19-03-S2, within the project “Fueling quantum field machines with information”). S.G. acknowledges support from the Knut and Alice Wallenberg foundation via the Wallenberg Centre for Quantum Technology (WACQT), from the European Research Council (Grant 101041744 ESQuAT), from the European Union (Grant 101080167 ASPECTS), and from the Swedish Research Council (Grant 2021-05624).
@@ Appendix A Non-Markovianity’s potential for reviving a machine’s purity
We follow [175]. That reference presents a setup common in cavity quantum electrodynamics: Consider a qubit $\mathcal{S}$ governed by a Hamiltonian $H_{\mathcal{S}}=\Delta\,\sigma_{z}$ with an energy splitting $\Delta>0$. As usual, we label the ground state as $\lvert 0\rangle$ and the excited state as $\lvert 1\rangle$. A bosonic environment $\mathcal{E}$ evolves under a Hamiltonian $H_{\mathcal{E}}=\sum_{\ell}\omega_{\ell}\,a_{\ell}^{\dagger}a_{\ell}$. Mode $\ell$ corresponds to energy $\omega_{\ell}$, to an annihilation operator $a_{\ell}$, and to a creation operator $a_{\ell}^{\dagger}$. The mode begins in its vacuum state. In the rotating-wave approximation, $S$ couples to $\mathcal{E}$ via the Jaynes–Cummings Hamiltonian $H_{\rm int}=\sum_{\ell}\big{(}g_{\ell}\,\sigma_{+}a_{\ell}+g_{\ell}^{*}\sigma_{-}a_{\ell}^{\dagger}\big{)}$. The qubit has raising and lowering operators $\sigma_{+}\coloneqq\lvert 1\rangle\!\langle 0\rvert$ and $\sigma_{-}\coloneqq\lvert 0\rangle\!\langle 1\rvert$, and the $g_{\ell}$’s denote coupling constants. Denote by $\rho_{\mathcal{S}/\mathcal{E}}(t)$ the time-$t$ reduced state of $\mathcal{S}/\mathcal{E}$ in the interaction picture. Let $\mathcal{S}$ and $\mathcal{E}$ begin in a product state $\rho_{\mathcal{S}}(0)\otimes\rho_{\mathcal{E}}(0)$, $\rho_{\mathcal{E}}(0)$ being the bath’s vacuum state.
@@
Figure 4 illustrates effects of a non-Markovian bath, whose coupling is strong: $\gamma_{0}=50\lambda/2$. Figure 4(a) shows the decoherence function vs. time. $G(t)$ oscillates, enabling the qubit’s energy coherences to revive repeatedly. The oscillations’ envelope decays eventually, however, demonstrating a limitation of non-Markovianity as a tool. Figure 4(b) shows the purity $\mathcal{P}\bm{(}\rho_{\mathcal{S}}(t)\bm{)}={\rm Tr}\bm{(}\rho_{\mathcal{S}}(t)^{2}\bm{)}$ of a qubit state initialized to $(\lvert 0\rangle+\lvert 1\rangle)/\sqrt{2}$. Calculated from Eqs. (A1), the purity oscillates until asymptoting. $\mathcal{P}$ asymptotes because $\mathcal{S}$ ends up in $\lvert 0\rangle$, as the environment began in its vacuum state and so sucks the energy out of $\mathcal{S}$. However, the oscillations demonstrate that the environment’s non-Markovianity partially returns $\mathcal{S}$ to its initial state repeatedly. Figure 5(b) depicts the same quantities—$G(t)$ and the purity—in the absence of non-Markovianity: The coupling $\gamma_{0}=0.4\lambda$ is weak. Neither function oscillates. Hence non-Markovianity can help baths achieve the energy criterion A without threatening the purity criterion F as much as thermal baths do, for a time.
@@
Figure 5: Coherence and purity at weak coupling: The setup is mostly as for Fig. 5. However, the coupling $\gamma_{0}=0.4\lambda$ is weak. Unlike in Fig. 5, $G(t)$ and $\mathcal{P}$ do not oscillate.

### v3

@@ Abstract
Controlled quantum machines have matured significantly. A natural next step is to increasingly grant them autonomy, freeing them from time-dependent external control. For example, autonomy could pare down the classical control wires that heat and decohere quantum circuits; and an autonomous quantum refrigerator recently reset a superconducting qubit to near its ground state, as is necessary before a computation. Which fundamental conditions are necessary for realizing useful autonomous quantum machines? Inspired by recent quantum thermodynamics and chemistry, we posit conditions analogous to DiVincenzo’s criteria for quantum computing. Furthermore, we illustrate the criteria with multiple autonomous quantum machines (refrigerators, circuits, clocks, etc.) and multiple candidate platforms (neutral atoms, molecules, superconducting qubits, etc.). Our criteria are intended to foment and guide the development of useful autonomous quantum machines.
@@
    Autonomous quantum machines of a third type could detect, transduce, and transmit energy. Nature has produced machines of this type. Photoisomers are molecules operative across nature and technologies [98, 99, 100, 101]. Examples include retinal, found in the protein rhodopsin in our eyes [102]. A photoisomer has one shape in thermal equilibrium at biological temperatures. Absorbing a photon—say, from the sun—enables the molecule to switch configurations. The switching can galvanize chemical reactions leading to the experience of sight. Photoisomers exhibit quantum phenomena including coherence relative to the energy eigenbasis, as well as nonadiabatic evolution. Beyond photoisomers, chemistry features other autonomous quantum machines: Photosynthetic complexes transform light into chemical energy [103], enzymes interconvert molecules with help from tunneling [104], etc.
@@ I Definitions
Before presenting our criteria, we specify meanings for machine, autonomous, and quantum. Alternative definitions may exist, as the terms are broad. Still, we believe our definitions to be reasonable. Furthermore, specifying them will clarify our criteria:
@@
    Autonomous: A machine is autonomous if its total microscopic Hamiltonian is not changed by any agent using the machine. An autonomous machine’s user performs no thermodynamic work on the machine during the machine’s operation. Work tends to be defined through changes in the system-of-interest Hamiltonian, in quantum thermodynamics [80].555 For example, an adenosine-triphosphate molecule provides chemical work via hydrolyzation: one of its phosphates splits off from the rest of the molecule, which becomes adenosine diphosphate. The cleaving of the bond releases energy and changes the molecule’s Hamiltonian.
@@
    Relatedly, this Perspective does not posit criteria under which autonomous quantum machines generally outperform all classical counterparts. Rather, the Perspective provides criteria under which autonomous quantum machines are useful. We believe this goal to be worthwhile, ambitious, and reasonable for the near future. Farther in the future, another Perspective may posit criteria under which autonomous quantum machines generally beat all classical competitors.
@@ A Access to useful energy
Other energy sources provide work indirectly. They are nonequilibrium systems that contain free energy (or a generalization thereof), which a machine can harvest to produce work. Often, such systems have gradients of temperature, chemical potential, or other generalized thermodynamic forces. The commonest example consists of two heat baths. One, at a temperature $T_{\mathcal{C}}$, is out of equilibrium with a heat bath at a temperature $T_{\mathcal{H}}>T_{\mathcal{C}}$. Heat flows between the baths. This heat is not work, so the temperature gradient does not provide work directly. However, a machine can siphon off a fraction of the current and transform it partially (in accordance with Carnot’s theorem) into work. Such machines are called absorption machines [74]. Heat baths of two types have fueled superconducting-qubit machines: resonators [122] and thermal fields propagating through microwave waveguides [75]. Despite providing energy, baths can threaten the purity of a machine’s state. Yet, as discussed under criterion F, non-Markovian baths, which retain memories, can revive lost purity.
@@ B Processing unit or target
A machine’s purpose can be to operate on a target. For instance, a refrigerator’s target is the system that undergoes cooling. An engine’s target may be a battery that stores the energy extracted by the engine. A drone’s target may be the package to be delivered.
@@ C Interactions among the machine’s components
Figure 2: Interactions among components: The quantum absorption refrigerator and clock evolve under similar Hamiltonians. However, the machines’ qudits have different energy gaps: the refrigerator’s $\Delta_{\mathcal{H}}<\Delta_{\mathcal{C}}$, whereas the clock’s $\Delta_{\mathcal{H}}>\Delta_{\mathcal{C}}$. Also, the number of the target’s energy gaps can differ between machines. The gaps, with the baths’ temperatures ($T_{\mathcal{H}}$ and $T_{\mathcal{C}}$), determine the direction in which energy flows.
@@
Effective three-body interactions are necessary for autonomous cooling [25]. They can be effected perturbatively with simultaneous two-body interactions [32, 34, 46, 75]. The exchange (1) occurs only if $\Delta_{\mathcal{H}}+\Delta_{\mathcal{T}}=\Delta_{\mathcal{C}}$—under a condition called strict, or microscopic, energy conservation [131, 132]. The thermal qubits’ gaps and temperatures bias the interaction (1) rightward. Ending in its ground state, $\mathcal{T}$ has undergone cooling by the refrigerator.
@@
Identity operators $\mathbbm{1}$ are implicitly tensored on wherever necessary for each operator to act on the full Hilbert space.
@@ D Timekeeping mechanism
Timekeeping devices include the switches introduced in the previous subsection; clocks, which tick regularly; and timers, which announce when a programmable amount of time has passed. Below, we review a simple autonomous quantum clock. Then, we discuss timekeeping mechanisms not formed from physical devices. Finally, we delineate three purposes of autonomous quantum machines’ timekeeping mechanisms.
@@ E Structural integrity
Yet a machine need not consist of physically separated components. A photoisomer functions as an autonomous quantum energy transmitter, as explained in the introduction. The molecule contains an autonomous quantum clock, according to [134]. Similarly, photosynthetic complexes [103] and enzymes [104] are autonomous machines modeled usefully with quantum theory. A molecule can therefore form an autonomous quantum machine. So can an atom, arguably.777 The atomic levels analyzed in [2, 3, 25] interact with heat baths whose frequencies are filtered. One might count the filters as parts of the machine. For example, three atomic energy levels can form an autonomous engine or refrigerator [2, 3, 25]. Not only natural particles, but also artificial devices can form self-contained autonomous quantum machines: Superconducting qubits cannot separate if printed on the same chip. Neither can the dopants that host qubits on solid-state surfaces; and neither can the gate electrodes that define quantum dots, being grown or printed on a semiconductor chip.
@@ G Output worth the input
Each above figure of merit measures an output’s quality. The efficiency and COP compare outputs with their input costs. Comprehensive figures of merit capture such comparisons [177]. Inputs can include energy, time, and control. We illustrate energy with a single-atom engine and a quantum refrigerator. Then, we distinguish two input costs: the fabrication of an autonomous quantum machine and the preparation of a fabricated machine for one trial. If the preparation is simple enough, and the machine executes enough trials, the total output can outweigh a large fabrication cost.
@@ I Mobility (optional)
    The signal must, given condition (1), have a sufficiently high probability of affecting $\mathcal{B}$. The excitation’s momentum should direct the signal toward $\mathcal{B}$, and no intervening medium should swallow the signal. Once the excitation arrives, $\mathcal{B}$ should have a high probability of absorbing it. Fermi’s golden rule can mediate this probability, governing the rate $\Gamma_{\text{i}\mapsto\text{f}}$ at which $\mathcal{B}$ jumps from a state $\lvert\rm i\rangle$ to any energy-$E_{\rm f}$ state $\lvert\rm f\rangle$: $\Gamma_{\text{i}\mapsto\text{f}}=\frac{2\pi}{\hbar}\,|\langle\text{f}\rvert H^{\prime}\lvert\text{i}\rangle|^{2}\,\mu(E_{\rm f})$. $H^{\prime}$ denotes the excitation-induced perturbation to $\mathcal{B}$’s Hamiltonian. $\mu(E)$ denotes the density of $\mathcal{B}$’s energy-$E$ states. If $\mathcal{B}$ absorbs the signal through a photodetector, the external efficiency quantifies how effectively $\mathcal{B}$ satisfies requirement (2). One can achieve a high efficiency upon matching $\mathcal{A}$’s and $\mathcal{B}$’s impedances. Impedance matching has enabled nearly perfect photon absorption by a cavity-QED detector [184, 185]. The scheme leveraged the detector’s $\Lambda$-type energy-level structure, albeit nonautonomously. Apart from impedance matching, realistic interconnects entail other engineering challenges, such as losses in the interconnects.
@@
In another example, a qubit and a bosonic mode undergo a dispersive interaction. Such interactions are common in cavity QED, the qubit manifesting as a natural or artificial atom. The dispersive interaction is effective (approximate), and we omit the derivation [187]. During it, one attributes to the qubit an effective gap $2\Delta$; to the mode an effective frequency $\omega$, a creation operator $a^{\dagger}$, and an annihilation operator $a$; and to the coupling a strength $\chi$. We set $\hbar=1$. The dispersive Hamiltonian is
@@
If the qubit is excited (in $\lvert 1\rangle$), it effectively adds $\chi$ to the mode’s frequency. Analogously, if the mode is occupied, it effectively adds $2\chi$ to the qubit’s gap. Each subsystem therefore affects the other’s Hamiltonian.
@@ III Outlook
    Usable energy—especially energy extracted from heat baths—trades off with purity, as discussed below criterion F. Four strategies suggest themselves: (i) Take advantage of the “sufficient” in the “sufficient purity” criterion: Identify when purity is necessary, and waste no effort on maintaining unnecessary purity. (ii) Situate autonomous quantum machines in environments that stabilize desirable quantum states [67], using engineered dissipation [189]. The engineered may suggest that maintaining the environment requires work from an agent. This impression is misleading, however. For example, “the Zeno effect [ …] is largely passive” and can raise a system’s probability of remaining in its ground state [189]. Engineered dissipation generalizes non-Markovian environments, discussed in App. A. (iii) Resolving challenge 1 can resolve this tradeoff challenge.
@@ Acknowledgements.
N.Y.H. thanks Kenneth Brown for input about castle-in-the-air autonomous quantum machines, David Limmer for input about chemical machines, Noah Lupu-Gladstein for input about photodetectors, Misha Lukin for input about Rydberg-atom operations, John Preskill for input about superconducting-qubit chips, and Nathan Schine for input about quantum-computing platforms. N.Y.H. and J.A.M.G. acknowledge support from the National Science Foundation (QLCI grant OMA-2120757), the John Templeton Foundation (award no. 62422), and NIST grant 70NANB21H055_0. P.E. and M.H. acknowledge support from the European Research Council (Consolidator grant “Co-coquest” 101043705), the European flagship on quantum technologies (“ASPECTS” consortium 101080167), and FQXi (FQXi-IAF19-03-S2, within the project “Fueling quantum field machines with information”). S.G. acknowledges support from the Knut and Alice Wallenberg foundation via the Wallenberg Centre for Quantum Technology (WACQT), from the European Research Council (Grant 101041744 ESQuAT), from the European Union (Grant 101080167 ASPECTS), and from the Swedish Research Council (Grant 2021-05624).
@@ Appendix A Non-Markovianity’s potential for reviving a machine’s purity
We follow [175]. That reference presents a setup common in cavity quantum electrodynamics: Consider a qubit $\mathcal{S}$ governed by a Hamiltonian $H_{\mathcal{S}}=\Delta\,\sigma_{z}$ with an energy splitting $02\Delta>$ 0. As usual, we label the ground state as $\lvert 0\rangle$ and the excited state as $\lvert 1\rangle$. A bosonic environment $\mathcal{E}$ evolves under a Hamiltonian $H_{\mathcal{E}}=\sum_{\ell}\omega_{\ell}\,a_{\ell}^{\dagger}a_{\ell}$. Mode $\ell$ corresponds to energy $\omega_{\ell}$, to an annihilation operator $a_{\ell}$, and to a creation operator $a_{\ell}^{\dagger}$. The mode begins in its vacuum state. In the rotating-wave approximation, $\mathcal{S}$ couples to $\mathcal{E}$ via the Jaynes–Cummings Hamiltonian $H_{\rm int}=\sum_{\ell}\big{(}g_{\ell}\,\sigma_{+}a_{\ell}+g_{\ell}^{*}\sigma_{-}a_{\ell}^{\dagger}\big{)}$. The qubit has raising and lowering operators $\sigma_{+}\coloneqq\lvert 1\rangle\!\langle 0\rvert$ and $\sigma_{-}\coloneqq\lvert 0\rangle\!\langle 1\rvert$, and the $g_{\ell}$’s denote coupling constants. Denote by $\rho_{\mathcal{S}/\mathcal{E}}(t)$ the time-$t$ reduced state of $\mathcal{S}/\mathcal{E}$ in the interaction picture. Let $\mathcal{S}$ and $\mathcal{E}$ begin in a product state $\rho_{\mathcal{S}}(0)\otimes\rho_{\mathcal{E}}(0)$, $\rho_{\mathcal{E}}(0)$ being the bath’s vacuum state.
@@
Figure 4 illustrates effects of a non-Markovian bath, whose coupling is strong: $\gamma_{0}=50\lambda/2$. Figure 4(a) shows the decoherence function vs. time. $G(t)$ oscillates, enabling the qubit’s energy coherences to revive repeatedly. The oscillations’ envelope decays eventually, however, demonstrating a limitation of non-Markovianity as a tool. Figure 4(b) shows the purity $\mathcal{P}\bm{(}\rho_{\mathcal{S}}(t)\bm{)}={\rm Tr}\bm{(}\rho_{\mathcal{S}}(t)^{2}\bm{)}$ of a qubit state initialized to $(\lvert 0\rangle+\lvert 1\rangle)/\sqrt{2}$. Calculated from Eqs. (A1), the purity oscillates until asymptoting. $\mathcal{P}$ asymptotes because $\mathcal{S}$ ends up in $\lvert 0\rangle$, as the environment began in its vacuum state and so sucks the energy out of $\mathcal{S}$. However, the oscillations demonstrate that the environment’s non-Markovianity partially returns $\mathcal{S}$ to its initial state repeatedly. Figure 5 depicts the same quantities—$G(t)$ and the purity—in the absence of non-Markovianity: The coupling $\gamma_{0}=0.4\lambda$ is weak. Neither function oscillates. Hence non-Markovianity can help baths achieve the energy criterion A without threatening the purity criterion F as much as thermal baths do, for a time.
@@
Figure 5: Coherence and purity at weak coupling: The setup is mostly as for Fig. 4(b). However, the coupling $\gamma_{0}=0.4\lambda$ is weak. Unlike in Fig. 4(b), $G(t)$ and $\mathcal{P}$ do not oscillate.